OpenStudy (anonymous):

sin (2sin^-1(x) PLEASE HELP BIG MATH TEST TOMRW

jimthompson5910 (jim_thompson5910):

I'm going to use arcsin(x) instead of sin^-1(x) Let y = arcsin(x) which means sin(y) = x/1 so sin(2*arcsin(x)) = sin(2y) Use the identity sin(2x) = 2*sin(x)*cos(x) to get sin(2y) = 2*sin(y)*cos(y) Now we turn to a drawing of a right triangle 

jimthompson5910 (jim_thompson5910):

cos(y) = adj/hyp = sqrt(1-x^2)/1 = sqrt(1-x^2)

jimthompson5910 (jim_thompson5910):

So... sin(2y) = 2*sin(y)*cos(y) sin(2y) = 2*x*sqrt(1-x^2) sin(2*arcsin(x)) = 2*x*sqrt(1-x^2)

OpenStudy (anonymous):

i seriously might fail tomorrow but thank you so much for your help

jimthompson5910 (jim_thompson5910):

is this all looking completely foreign to you?

OpenStudy (anonymous):

no i'm really good with trig identities, i'm just lost when solving equations. i solved this question right but it looked wrong so i thought i needed help

jimthompson5910 (jim_thompson5910):

so the identity sin(2y) = 2*sin(y)*cos(y) looks familiar?

OpenStudy (anonymous):

yes thank you

jimthompson5910 (jim_thompson5910):

So what you need to do is go from sin(2*arcsin(x)) to sin(2y) Then break that up into 2*sin(y)*cos(y)

jimthompson5910 (jim_thompson5910):

then use the diagram to find sin(y) and cos(y) sin(y) is already given based on how you defined y since y = arcsin(x), this means sin(y) = x